Two ice skaters suddenly push off against one another starting from a stationary position. The 45-kg skater acquires a speed of 0.375 m/s relative to the ice. What speed does the 60-kg skater acquire relative to the ice?


1
Expert's answer
2021-02-25T11:39:22-0500

Answer

Momentum is conserved in this case.

Initial momentum = Final momentum.

Initial momentum = Initial momentum of skater 1 + Initial momentum of skater 2

Initial momentum

"= 45 \times0 + 60 \times 0 \\= 0 kg m/s"

Final momentum

"= 45 \times 0.375 + 60 \times u \\= 16.875 + 60 u"

Here we need to find u.

We have

        "0 = 16.875 + 60 u\\

 u = -0.281 m/s"

Speed does the 60-kg skater acquire relative to the ice"= 0.281 m\/s"